OpenStudy (anonymous):

I WILL GIVE MEDALS What is the simplified form of (x^(2)-25)/(x-5) ? x - 5, with the restriction x ≠ -5 x + 5, with the restriction x ≠ 5 x - 5, with the restriction x ≠ 5 x + 5, with the restriction x ≠ -5

OpenStudy (inkyvoyd):

number 2

OpenStudy (inkyvoyd):

x cannot be 5, because that's undefined. difference of squares: (a+b)(a-b)=a^2-b^2

OpenStudy (campbell_st):

the numerator factorises to(x -5)(x+5) eliminate the common factor to get the answer

OpenStudy (campbell_st):

restriction is what value of x make the denominator zero x - 5 = 0